Five dollars is divided continuously among three boys the first boy always gets $.02 the second boy always gets $.03 and the third always gets $.05 until all the money is gone how much money does each boy gets?
Answer by ikleyn(52802)

.

It means that the boys will get money in proportion 2 : 3 : 5.


In other words, 1-st boy will get 2x dollars; 

                2-nd boy will get 3x dollars,

            and 3-rd boy will get 5x dollars.


To find x, use the total money equation

    2x + 3x + 5x = 5   dollars

    10x          = 5

      x          = 5/10 = 0.5 dollars.


Thus 1-st boy will get 2*0.5 = 1 dollar; 2-nd will get 3*0.5 = 1.5 dollards, and 3-rd will get 5*0.5 = 2.5 dollars.   ANSWER

Solved.